# Random Turán Problems for Hypergraph Expansions

### v1

@@ A.1 Comparing the Methods
Here we very informally show that 1.7, whose proof used both Greedy and Shadow Expansion, is essentially always at least as strong as 1.6, whose proof only used Shadow Expansion. For this we use the notation $O^{*}$ and $\Omega^{*}$ to denote asymptotic upper bounds and lower bounds that hold up to polylogarithmic factors, and throughout our argument we will be loose with logarithmic terms in order to make some of the calculations clearer.
Let $F$ be an $(r-1)$-graph which satisfies the hypothesis of 1.7 by being $(M_{r-1},\gamma_{r-1},\tau_{r-1})$-balanced, and let us make the extra mild assumption that $\gamma_{r-1}(n)=O^{*}(1)$. As we show in 3.1, this implies $\tau_{r-1}(n,m)=\Omega^{*}(n^{r-\frac{1}{d_{r}(F^{(r)})}})$. With this in mind, we consider applying 1.7 with $A=(m/n)^{1/(r-1)}$ for $m\geq\max\{M_{r-1}^{\frac{r-1}{r-2}}(n)n^{-\frac{1}{r-2}},n^{r-1}\}$ (i.e. for the range of $m$ that 1.6 is valid for). Note that $m\geq M_{r-1}^{\frac{r-1}{r-2}}(n)n^{-\frac{1}{r-2}}$ is equivalent to having $A=(m/n)^{1/(r-1)}\leq m/M_{r-1}(n)$, so (up to logarithmic factors) we have that the upper bound of property (d) in 1.7, and it is straightforward to verify the lower bound $A\geq mn^{1-r}$ for $r\geq 3$ and $m\geq n^{r-1}$, so we can indeed apply 1.7 with this value of $A$. We now observe that the first term in the maximum of $\tau_{r}(n,m)$ defined in 1.7 satisfies
@@
with the inequality using $m\leq n^{r}$. This together with the claim above implies that up to logarithmic terms $\tau_{r}(n,m)$ equals $\tau_{r-1}(n,A^{-1}m)=\tau_{r-1}(n,n^{\frac{1}{r-1}}m^{\frac{r-2}{r-1}})$, and this is exactly what 1.6 gives.
A.2 Expansions of Complete Bipartite Graphs
Here we prove 1.3 giving tight bounds for random Turán numbers of $K_{s,t}$ expansions for $r$ sufficiently large. We need the following lemma.
@@ A.2 Expansions of Complete Bipartite Graphs
By 3.3 it suffices to construct a tight $(s+1)$-tree whose 2-shadow contains a copy of $K_{s,t}$. Let $S$ and $T$ be the two parts of $K_{s,t}$ such that $|S|=s$ and $|T|=t$, where $S=\{v_{1},\dots,v_{s}\}$ and $T=\{w_{1},\dots,w_{t}\}$. Pick the $(s+1)$-edges $e_{i}=S\cup\{w_{i}\}$ for $1\leq i\leq t$. It is easy to check that $e_{1},\dots,e_{t}$ form a tight $(s+1)$-tree whose 2-shadow contains all $v_{i}w_{j}$, where $1\leq i\leq s$ and $1\leq j\leq t$. ∎
@@
We will first use 1.6 to obtain balanced supersaturation results for $K^{(s)}_{s,t}$, and then apply 1.7 to deal with $K^{(r)}_{s,t}$ when $r\geq s+1$.
Let $M_{2},C_{2},\tau_{2}$ be as in Theorem 5.3 and let $\gamma_{2}(n)$ be constantly $C_{2}$. In order to apply 1.6 with $F=K_{s,t}$ and $r=s$, we need to check the four properties (a), (b), (c) and (d). Property (a) of $F$ being $(M_{2},\gamma_{2},\tau_{2})$-balanced is immediate from Theorem 5.3, property (b) of having $M_{2}(n)n^{-\frac{s-2}{s-1}}=n^{1+\frac{1}{s(s-1)}}$ being non-increasing holds, and property (c) of $\gamma_{2}(n)$ being non-decreasing is immediate. For property (d), we observe that $\tau_{2}(n,m)$ is non-increasing with $m$ and that for every $s\geq 2$
@@
is increasing with respect to $x$. So property (d) also holds. Hence by 1.6, there exists sufficiently large $C_{s}$ such that for the functions
@@
We use Lemma A.2 together with 1.5 to prove Theorem 1.3.
@@
By 4.1 it suffices to only prove the upper bound of Theorem 1.3 for the larger range of $p$. By Lemma A.2, we can apply Lemma 1.5 with $F=K^{(r)}_{s,t}$ which guarantees the existence of a sufficiently large constant $C_{1}$ such that for all $n\geq N_{r}$ and $m\geq C_{r}n^{r-1}$,

### v2

@@ A.1 Comparing the Methods
Here we very informally show that 2.8, whose proof used both Greedy and Shadow Expansion, is essentially always at least as strong as 2.7, whose proof only used Shadow Expansion. For this we use the notation $O^{*}$ and $\Omega^{*}$ to denote asymptotic upper bounds and lower bounds that hold up to polylogarithmic factors, and throughout our argument we will be loose with logarithmic terms in order to make some of the calculations clearer.
Let $F$ be an $(r-1)$-graph which satisfies the hypothesis of 2.8 by being $(M_{r-1},\gamma_{r-1},\tau_{r-1})$-balanced, and let us make the extra mild assumption that $\gamma_{r-1}(n)=O^{*}(1)$. As we show in 4.1, this implies $\tau_{r-1}(n,m)=\Omega^{*}(n^{r-\frac{1}{d_{r}(F^{(r)})}})$. With this in mind, we consider applying 2.8 with $A=(m/n)^{1/(r-1)}$ for $m\geq\max\{M_{r-1}^{\frac{r-1}{r-2}}(n)n^{-\frac{1}{r-2}},n^{r-1}\}$ (i.e. for the range of $m$ that 2.7 is valid for). Note that $m\geq M_{r-1}^{\frac{r-1}{r-2}}(n)n^{-\frac{1}{r-2}}$ is equivalent to having $A=(m/n)^{1/(r-1)}\leq m/M_{r-1}(n)$, so (up to logarithmic factors) we have that the upper bound of property (d) in 2.8, and it is straightforward to verify the lower bound $A\geq mn^{1-r}$ for $r\geq 3$ and $m\geq n^{r-1}$, so we can indeed apply 2.8 with this value of $A$. We now observe that the first term in the maximum of $\tau_{r}(n,m)$ defined in 2.8 satisfies
@@
with the inequality using $m\leq n^{r}$. This together with the claim above implies that up to logarithmic terms $\tau_{r}(n,m)$ equals $\tau_{r-1}(n,A^{-1}m)=\tau_{r-1}(n,n^{\frac{1}{r-1}}m^{\frac{r-2}{r-1}})$, and this is exactly what 2.7 gives.
A.2 Proof of 6.2
We make use of the following lemma of Jiang and Longbrake [13].
Lemma A.1 (Lemma 3.4, [13]).
Let $r,h,\ell$ be positive integers, where $h\geq r\geq 2$. Let $F$ be an $r$-partite $r$-graph with $h$ vertices and $\ell$ edges. Let $\alpha,C$ be positive reals such that for each $n$ every $n$-vertex $r$-graph $H$ with $m\geq Cn^{\alpha}$ edges contains at least $f(n,m)$ copies of $H$, where $f$ is a function satisfying the following.
  • (a)
    There is a constant $\delta>0$ such that for all $p\in(0,1]$ and all positive reals $n,m$
    $f(n,m)\geq\delta m^{\frac{h-\alpha}{r-\alpha}}/n^{\frac{\alpha(h-r)}{r-\alpha}}\quad\mbox{ and }\quad f(np,mp^{r})\geq\delta f(n,m)p^{h}.$
  • (b)
    For fixed $n$, $f(n,m)$ is increasing and convex in $m$.
Then there exist a constant $C^{\prime}=C^{\prime}(F)\geq 1$ such that if $H$ is an $n$-vertex $r$-graph with $kn^{\alpha}$ edges, where $k\geq C^{\prime}$, then there exists a family $\mathcal{F}$ of copies of $F$ in $H$ satisfying
  • (c)
    $|\mathcal{F}|\geq\delta f(n,\frac{1}{8}|H|)$, and
  • (d)
    $\forall 1\leq i\leq\ell-1$,
    $\Delta_{i}(\mathcal{F})\leq C^{\prime}k^{-\frac{i-1}{d^{*}_{r}(F)(r-\alpha)}}\frac{|\mathcal{F}|}{|H|},$
    where $d^{*}_{r}(F)$ is the proper $r$-density of $F$.
We are now ready to prove 6.2, which we restate below for convenience.
Theorem 6.2.
Let $F$ be a non-empty $r$-partite $r$-graph with $|F|\geq 2$. If there exist positive reals $C$ and $\alpha>r-\frac{1}{d_{r}(F)}$ such that $F$ is $(C,\alpha)$-Sidorenko, then there exists constant $C^{\prime}=C^{\prime}(F)>0$ such that $F$ is $(M,\gamma,\tau)$-balanced where $M(n)=C^{\prime}n^{r-\frac{1}{|F|}}$, $\gamma=C^{\prime}$, and
$\tau(n,m)=C^{\prime}\max\left\{n^{r-\frac{v(F)-r}{|F|-1}},~{}m^{1-\frac{1}{d_{r}^{*}(F)(r-\alpha)}}n^{\frac{\alpha}{d^{*}_{r}(F)(r-\alpha)}}\right\}.$
Proof.
Since $F$ is $(C,\alpha)$-Sidorenko, for every $n$-vertex $r$-graph $H$ with $m\geq Cn^{\alpha}$ edges contains at least $C^{-1}m^{|F|}n^{-r|F|+v(F)}$ copies of $F$. We want to apply A.1 with
$f(n,m)=C^{-1}m^{|F|}n^{-r|F|+v(F)}.$
To this end, we need to check that conditions (a) and (b) in A.1 are satisfied. Since $|F|\geq 2$, we know that $v(F)\geq r+1$, and hence $m^{\frac{v(F)-\alpha}{r-\alpha}}/n^{\frac{\alpha(v(F)-r)}{r-\alpha}}$ is maximized when $\alpha$ is minimized. Thus by $\alpha>r-\frac{1}{d_{r}(F)}$ and $m\leq n^{r}$ we have
$\frac{f(n,m)n^{\frac{\alpha(v(F)-r)}{r-\alpha}}}{m^{\frac{v(F)-\alpha}{r-\alpha}}}>C^{-1}m^{|F|-1-d_{r}(v(F)-r)}n^{-r|F|+v(F)+(d_{r}r-1)(v(F)-r)}=C^{-1}\left(\frac{n^{r}}{m}\right)^{d_{r}(v(F)-r)-|F|+1}\geq C^{-1}.$
Also note that
$\frac{f(np,mp^{r})}{f(n,m)p^{v(F)}}=1.$
Thus condition (a) in A.1 holds with $\delta=C^{-1}$. Further, note that the exponent of $m$ in $f(n,m)$ is $|F|\geq 2$, thus condition (b) holds. By A.1, there exists a constant $C_{1}\geq 1$ such that if $H$ is an $n$-vertex $r$-graph with $m\geq C_{1}n^{\alpha}$ edges, then there exists a family $\mathcal{F}$ of copies of $F$ in $H$ such that
$|\mathcal{F}|\geq f(n,\frac{1}{8}m)=C^{-1}8^{-|F|}m^{|F|}n^{-r|F|+v(F)},$ (22)
and, $\forall 1\leq i\leq|F|-1$,
$\Delta_{i}(\mathcal{F})\leq C_{1}\left({m}{n^{-\alpha}}\right)^{-\frac{i-1}{d^{*}_{r}(F)(r-\alpha)}}\frac{|\mathcal{F}|}{|H|}=O\left(\left(\frac{\tau(n,m)}{m}\right)^{i-1}\frac{|\mathcal{F}|}{|H|}\right),$
where the last equal sign comes from using the second term in the maximum in $\tau(n,m)$.
Note that by Equation 22,
$1=\Delta_{|F|}(\mathcal{F})\leq C_{1}8^{|F|}m^{-|F|+1}n^{r|F|-v(F)}\frac{|\mathcal{F}|}{|H|}=O\left(\left(\frac{\tau(n,m)}{m}\right)^{|F|-1}\frac{|\mathcal{F}|}{|H|}\right),$
where the last equal sign comes from using the first term in the maximum in $\tau(n,m)$. Therefore, the proof is completed by taking sufficiently large $C^{\prime}$. ∎
A.3 Expansions of Complete Bipartite Graphs
Here we prove 2.3 giving tight bounds for random Turán numbers of $K_{s,t}$ expansions for $r$ sufficiently large. We need the following lemma.
@@ A.3 Expansions of Complete Bipartite Graphs
By 4.3 it suffices to construct a tight $(s+1)$-tree whose 2-shadow contains a copy of $K_{s,t}$. Let $S$ and $T$ be the two parts of $K_{s,t}$ such that $|S|=s$ and $|T|=t$, where $S=\{v_{1},\dots,v_{s}\}$ and $T=\{w_{1},\dots,w_{t}\}$. Pick the $(s+1)$-edges $e_{i}=S\cup\{w_{i}\}$ for $1\leq i\leq t$. It is easy to check that $e_{1},\dots,e_{t}$ form a tight $(s+1)$-tree whose 2-shadow contains all $v_{i}w_{j}$, where $1\leq i\leq s$ and $1\leq j\leq t$. ∎
@@
We will first use 2.7 to obtain balanced supersaturation results for $K^{(s)}_{s,t}$, and then apply 2.8 to deal with $K^{(r)}_{s,t}$ when $r\geq s+1$.
Let $M_{2},C_{2},\tau_{2}$ be as in Theorem 6.8 and let $\gamma_{2}(n)$ be constantly $C_{2}$. In order to apply 2.7 with $F=K_{s,t}$ and $r=s$, we need to check the four properties (a), (b), (c) and (d). Property (a) of $F$ being $(M_{2},\gamma_{2},\tau_{2})$-balanced is immediate from Theorem 6.8, property (b) of having $M_{2}(n)n^{-\frac{s-2}{s-1}}=n^{1+\frac{1}{s(s-1)}}$ being non-increasing holds, and property (c) of $\gamma_{2}(n)$ being non-decreasing is immediate. For property (d), we observe that $\tau_{2}(n,m)$ is non-increasing with $m$ and that for every $s\geq 2$
@@
is increasing with respect to $x$. So property (d) also holds. Hence by 2.7, there exists sufficiently large $C_{s}$ such that for the functions
@@
We use Lemma A.3 together with 2.6 to prove Theorem 2.3.
@@
By 2.5 it suffices to only prove the upper bound of Theorem 2.3 for the larger range of $p$. By Lemma A.3, we can apply Lemma 2.6 with $F=K^{(r)}_{s,t}$ which guarantees the existence of a sufficiently large constant $C_{1}$ such that for all $n\geq N_{r}$ and $m\geq C_{r}n^{r-1}$,